### v1

# Impurity dynamics in a zero-temperature gas

Abstract
If energy is suddenly released in a localized region of space uniformly filled with identical stationary hard spheres of radii $a$, the outcome is a blast with an asymptotically spherical shock wave separating moving and stationary hard spheres. The radius of the region filled with the moving spheres grows as $R\sim(Et^{2}/n_{\infty})^{1/(d+2)}$, where $d$ is the spatial dimension, $n_{\infty}$ the number density of stationary spheres and $E$ is the total energy injected. The simplest way to inject energy is to kick a few ‘impurity’ particles. Using hydrodynamics and kinetic theory, we argue that the typical displacement of an impurity scales as $R_{\rm imp}\sim\lambda(R/\lambda)^{(4+3d^{2})/(8+3d^{2})}$, where $\lambda\sim 1/(n_{\infty}a^{d-1})$ is the mean-free path in the initial state. The number of collisions experienced by each impurity is $(R/\lambda)^{(8+2d^{2})/(8+3d^{2})}$, while its average speed decreases as $t^{-d(8-2d+3d^{2})/[(2+d)(8+3d^{2})]}$. In $2D$, the predictions for impurity displacement, collision numbers, and speed are $t^{2/5},~{}t^{2/5}$ and $t^{-2/5}$, respectively. These predictions are in reasonable agreement with the results of molecular dynamics simulations.
I Introduction
A localized instantaneous release of energy generates a blast, first investigated in the context of atomic explosions Taylor (1950a, b); Sedov (1959); von Neumann (1963). The blast is strong as long as the pressure behind the shock wave greatly exceeds the atmospheric pressure. For the strong blast, Euler equations describing an ideal compressible gas admits an exact solution; see textbooks Zeldovich and Raiser (1967); Landau and Lifshitz (1987); Barenblatt (1996). The shock wave eventually weakens, and the problem becomes analytically intractable. However, we look at the idealized case where the pressure in front of the shock is zero, so the blast is infinitely strong throughout the evolution. Infinitely strong blasts mimic supernova explosions and other astrophysical phenomena Ostriker and McKee (1988); Tang and Chevalier (2017); Vink (2020); Modelevsky and Sari (2021).
One can examine the infinitely strong blast problem at a microscopic level by considering a hard sphere gas with particles initially at rest, i.e., at zero temperature. How does this system evolve when energy is injected into a localized region, for example, by giving energy to a few particles? An infinite system with initially immobile particles has an additional virtue—it is amenable to direct molecular dynamics simulations since there is only a finite number of moving particles. Hence, one can investigate the behavior in a numerically precise way. This system thus provides a fruitful laboratory for probing the validity of hydrodynamics, explaining the popularity of the setting with initially immobile hard spheres and the number of recent papers studying the blast for this system Antal et al. (2008); Barbier et al. (2015, 2016); Joy et al. (2021); Joy and Rajesh (2021); Kumar and Rajesh (2022); Chakraborti et al. (2021); Ganapa et al. (2021); Chakraborti et al. (2022); Kumar and Rajesh (2024, 2025a); Singh et al. (2023). These studies compared the predictions of Euler hydrodynamics with the results from microscopic simulations. Somewhat surprisingly, one finds rather good agreement. However, deviations from the predictions of the Euler equations have been observed. These deviations can be understood by taking into account dissipative effects such as heat conduction in a core region around the center of the blast, i.e., one needs to use the full Navier-Stokes equations Chakraborti et al. (2021); Ganapa et al. (2021); Singh et al. (2023).
Previous studies focused on coarse-grained observables such as density, temperature, and pressure. The emergence of the hydrodynamic behavior is questionable, at least at first sight, as the gas outside the shock wave has zero temperature. The success of hydrodynamics comes as a pleasant surprise. Going beyond the scope of conventional hydrodynamics predicting macroscopic characteristics, one can study mesoscopic characteristics such as the self-diffusion phenomenon inside a blast. The present work explores precisely this question. Employing hydrodynamics and kinetic theory arguments we deduce scaling laws describing the characteristics of the impurity.
[…]
In Sec. II, we describe the model and summarize the main results. In Sec. III, we derive the main results using arguments based on kinetic theory. In Sec. IV, we argue that the distributions $P(\bm{r},t)$ and $Q(\bm{v},t)$ of the position and velocity of the impurity do not satisfy closed equations. The joint position-velocity distribution $\Pi(\bm{r},\bm{v},t)$ provides a more comprehensive description than the individual distributions $P(\bm{r},t)$ and $Q(\bm{v},t)$ together. The joint position-velocity distribution $\Pi(\bm{r},\bm{v},t)$ satisfies a closed Lorentz-Boltzmann equation. The integro-differential Lorentz-Boltzmann equation in an inhomogeneous and evolving background seems analytically intractable. In Sec. V, we present the results of simulations and their comparison with the analytic predictions. We conclude with a summary in Sec. VI.
II The model and summary of results
Consider a homogeneous hard-sphere gas with initially static spheres, each of radius $a$, and unit mass $m=1$. We perturb the system at time $t=0$ by giving energy, $E$, to a few particles Antal et al. (2008); Krapivsky et al. (2010). We will refer to these particles as the impurities. The impurities elastically collide with the host spheres, which begin to move and elastically collide with other spheres, generating a growing number of moving spheres. We are primarily interested in the ‘fate’ of the impurities, specifically the asymptotic distributions of their position and velocity. In Fig. 1, we present a simulation of a blast in a two-dimensional hard-disc gas initiated by injecting energy into four impurity particles. The figure also shows the trajectories of these impurity particles against the backdrop of the evolving blast environment.
Following the classic works of Taylor-von-Neumann-Sedov (TvNS) Taylor (1950a, b); von Neumann (1963); Sedov (1959) on the self-similar solution of the Euler equations, it is known that the moving spheres are separated from quiescent spheres by an asymptotically spherical shock wave advancing as
$R=\left(\frac{Et^{2}}{A_{d}n_{\infty}}\right)^{\frac{1}{d+2}}$ (1)
where $n_{\infty}$ is the number density of spheres and $d$ is the spatial dimension. At first sight, the sub-linear growth $R\sim t^{2/(d+2)}$ is surprising because the shock wave always moves faster than the sonic speed Landau and Lifshitz (1987). However, the sonic speed vanishes in the zero-temperature gas, so, albeit the speed of the shock decays as $t^{-d/(d+2)}$, the shock remains infinitely strong throughout the evolution. The position of the shock wave depends only on macroscopic variables, i.e., total energy, $E$, and number density, $n_{\infty}$. The notation $A_{d}$ for the amplitude in Eq. (1) emphasizes the dependence on the spatial dimension, but $A_{d}$ also depends on the volume fraction
$\phi=n_{\infty}\,\Omega_{d}a^{d},\quad\Omega_{d}=\frac{\pi^{\frac{d}{2}}}{\Gamma\big{(}\frac{d}{2}+1\big{)}}$ (2)
occupied by spheres. These amplitudes are generally unknown since even the equation of state for the dense hard-sphere gas is unknown. In the $\phi\to 0$ limit, the hard-sphere gas is dilute and ideal. Euler equations for ideal compressible gas behind the shock arising in our infinitely strong blast problem admit an exact solution for arbitrary adiabatic index $\gamma$. We set $\gamma=1+2/d$, the adiabatic index characterizing ideal monoatomic gases Resibois and Leener (1977); Dorfman et al. (2021); Lifshitz and Pitaevskii (1981). In the physically relevant dimensions, the amplitudes are
[…]
However, the hard-sphere simulations revealed deviations from the exact solution of the Euler equations for the ideal compressible gas. One should account for heat conduction in the core region which grows slower than the shock wave, i.e., rely on the Navier-Stokes equations Chakraborti et al. (2021); Ganapa et al. (2021); Singh et al. (2023). There is no sharp boundary separating the core from the bulk described by Euler equations, but one can still talk about the characteristic radius of the core region. The temperature in the core decays with time, yet the ratio of this temperature to the temperature in the bulk of the moving gas behind the shock diverges with time. Thus, the core region is ‘hot’; we denote its characteristic radius by $H$. This radius grows as Singh et al. (2023)
$H=\lambda\left(\frac{R}{\lambda}\right)^{h_{d}},$ (3)
and we revisit the derivation of this formula in the next section.
Our primary interest in the present work is the ‘fate’ of the impurities — specifically the asymptotic distributions of their position, velocity, and other relevant quantities. We summarize below our main findings.
  • As opposed to the position of the shock front, which becomes asymptotically deterministic, the position of the impurity remains a non-self-averaging random quantity. We argue that in the long-time limit, the typical impurity displacement depends on the shock front radius $R$ and the mean-free path in the initial state $\lambda$ as:
    $R_{\text{imp}}=\lambda\left(\frac{R}{\lambda}\right)^{h_{d}},\qquad h_{d}=\frac{4+3d^{2}}{8+3d^{2}},$ (4)
    where the mean-free path is given by
    $\lambda=\frac{1}{n_{\infty}\Omega_{d-1}a^{d-1}}.$ (5)
    In two dimensions, this yields $R_{\text{imp}}\propto t^{2/5}$. We test this prediction in simulations and a comparison with the theory is shown in Fig. 2(A).
  • The typical value of the impurity velocity is given by
    $V_{\text{imp}}=V\left(\frac{a}{\lambda}\right)^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\omega_{d}},\quad\omega_{d}=\frac{d}{2}-\frac{d^{2}}{8+3d^{2}}.$ (6)
    Since ${a}/{\lambda}\sim\phi$, the volume fraction (see (2) and (5)), we can re-write (6) as
    $V_{\text{imp}}=V\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\omega_{d}}.$ (7)
    For the two-dimensional case, this gives $V_{\text{imp}}\propto t^{-2/5}$ and a comparison between theory and simulation is shown in Fig. 2(B). On the other hand, the hydrodynamic flow near the impurity is predominantly radial and follows the relation $u_{\text{imp}}\sim R_{\text{imp}}/t$. More precisely, we find
    $u_{\text{imp}}=V\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\frac{d}{2}+1-h_{d}}.$ (8)
    For two dimensions, this gives $u_{\text{imp}}\propto t^{-3/5}$. A comparison between theory and simulation is shown in Fig. 2(D).
  • We show that the typical number of collisions, $C_{\rm imp}$, experienced by the impurity scales as
    $C_{\text{imp}}=\left(\frac{R}{\lambda}\right)^{\frac{8+2d^{2}}{8+3d^{2}}}.$ (9)
    For the two-dimensional case, this leads to $C_{\text{imp}}\propto t^{2/5}$. A comparison between theory and simulation is shown in Fig. 2(C).
  • As part of our heurisitic derivation of some of the above results, we also show that in the core region, the density $n_{0}$ and temperature $T_{0}$ are given by
    $n_{0}=n_{\infty}\left(\frac{\lambda}{R}\right)^{d-2\omega_{d}}$ (10)
    and
    $T_{0}=mV^{2}\left(\frac{a}{\lambda}\right)^{d-1}\left(\frac{\lambda}{R}\right)^{2\omega_{d}}.$ (11)
    In two dimensions, $n_{0}\propto t^{-1/5}$ and $T_{0}\propto t^{-4/5}$; see Fig. 2(E)-(F) for the comparison with simulation results.
The mean free path depends on the radius of spheres, a microscopic variable. Therefore, in contrast to the position of the shock wave, (1), predicted by ideal compressible hydrodynamics, the determination of the displacement of the impurity seemingly requires a microscopic framework. Indeed, the derivation of (4) relies on the arguments used in kinetic theory Resibois and Leener (1977); Dorfman et al. (2021). Remarkably, the length scale $R_{\text{imp}}$ is identical to the one mentioned in (3), that arises in a more accurate hydrodynamic description of the flow in the core of the blast. Thus, the impurity remains inside the core.
The above formulas apply when $d\geq 2$, particularly in two and three dimensions. The one-dimensional setting is pathological due to the lack of hydrodynamic behavior: There is no mixing since two elastically colliding particles exchange the velocities and effectively pass through each other. In our setting, there is only one moving particle at any moment. One circumvents this pathology by allowing for distributed particle masses. The hydrodynamic behavior then emerges Antal et al. (2008); Chakraborti et al. (2021); Ganapa et al. (2021). The core region was originally detected in one dimension Chakraborti et al. (2021); Ganapa et al. (2021). The behavior of the impurity remains pathological since, in one dimension, the impurity is forever caged between the right and left neighbors. This caging property can be circumvented by postulating that the impurity colliding with a host particle passes with a certain probability $p$, while elastic collision occurs with probability $1-p$. (This trick allows one to devise a meaningful version of the one-dimensional Lorentz gas Piasecki and Wajnryb (1979); Krapivsky and Redner (1997).) With such an amendment, the displacement of the impurity should be comparable with the size of the core region. Below, we always assume that $d\geq 2$ and employ the basic setting with spheres of equal masses undergoing elastic collisions.
Figure 2: Time evolution of various impurity and fluid properties plotted on log-log scale: (A) impurity displacement $R_{\text{imp}}$; (B) impurity speed $V_{\text{imp}}$; (C) impurity collision count $C_{\text{imp}}$; (D) fluid speed around the impurity $u_{\text{imp}}$; (E) central density $n_{0}$; (F) central temperature $T_{0}$. All quantities are asymptotically expected to have power-law time dependence of the form $t^{\alpha}$ and the insets show the running time slopes plotted with time. The dark dashed lines correspond to the theoretical predictions.
III Heuristic Derivations
The strategy of our heuristic derivation is that we first establish the properties in the core region where dissipation (viscosity and thermal conductivity) is important. In particular, we find the typical size (3) of the core region, and the density (10) and temperature (11) in this region. Next, to justify the scaling laws (4) and (7) –(9), we assume that in the long time limit, the impurity is in the core region and demonstrate self-consistency thus confirming a posteriori that the impurity is inside the core. The derivation of Eqs. (4) and (7) –(9) then follows from arguments based on elementary kinetic theory.
The growth law $H\propto R^{h_{d}}$ for the characteristic size of the core region was established in Singh et al. (2023). We now derive the more accurate result (3). (Below, we also use the same approach to generalize (3) to monatomic gases with point particles and algebraically diverging repulsive interactions.) To estimate the size of the core region where dissipative effects matter, it suffices to keep heat conduction and ignore viscous dissipation Singh et al. (2023); Chakraborti et al. (2021); Ganapa et al. (2021). In this situation, the entropy equation reads
$\rho T(\partial_{t}+u\partial_{r})s=r^{-(d-1)}\,\partial_{r}(r^{d-1}\kappa\partial_{r}T),$ (12)
where $u$ is the hydrodynamic radial velocity,
$s=\frac{d}{2}\,\ln\!\left(\frac{T}{\rho^{2/d}}\right)$ (13)
the entropy per unit mass (assuming the ideal gas form), and $\kappa$ the coefficient of thermal conductivity. According to kinetic theory Resibois and Leener (1977); Dorfman et al. (2021); Lifshitz and Pitaevskii (1981), for the dilute hard-sphere gas, the coefficient of thermal conductivity is proportional to the thermal velocity $\sqrt{T/m}$ and inversely proportional to the cross-section area $\sigma\sim a^{d-1}$. It is convenient to write $\kappa$ as
$\kappa=n_{\infty}\lambda\sqrt{T}$ (14)
where we ignore (analytically unknown) numerical factor. Using (13) we recast (12) into
$\rho^{1+2/d}\,(\partial_{t}+v\partial_{r})\frac{T}{\rho^{2/d}}=\frac{1}{r^{d-1}}\,\partial_{r}\big{(}r^{d-1}\kappa\,\partial_{r}T\big{)}$ (15)
Inserting (14) into (15) we estimate the radius $H$ of the core where the heat transfer is comparable with the terms on the left-hand side
$\rho\sim n_{\infty}\lambda\,\sqrt{T}\,\frac{t}{H^{2}}$ (16)
From the exact solution outside the core Singh et al. (2023) one finds the asymptotic behaviors valid when $r\ll R$:
$\frac{\rho}{n_{\infty}}$ $=\left(\frac{r}{R}\right)^{\frac{d^{2}}{2}}$ (17a)
$T$ $=\left(\frac{R}{t}\right)^{2}\left(\frac{r}{R}\right)^{-\frac{d^{2}}{2}}$ (17b)
The divergence of the temperature when $r\to 0$ is physically impossible. The asymptotic behaviors (17a)–(17b) are valid for $r>H$. Setting $r=H$ in (17a)–(17b) and inserting into (16) we arrive at (3).
The density and temperature in the core region are comparable to the density $n_{0}$ and the temperature $T_{0}$ at the center of the explosion. The density $n_{0}$ is estimated by setting $r=H$ in (17a):
$n_{0}=n_{\infty}\left(\frac{\lambda}{R}\right)^{d-2\omega_{d}}$ (18)
The temperature $T_{0}$ is similarly determined by setting $r=H$ in (17b):
$T_{0}=V^{2}\left(\frac{a}{\lambda}\right)^{d-1}\,\left(\frac{\lambda}{R}\right)^{2\omega_{d}}$ (19)
The typical speed of the impurity is essentially the thermal speed $\sqrt{T_{0}}$ in the core. Using $V_{\text{imp}}=\sqrt{T_{0}}$ and (19) we arrive at (6) and (7).
To derive the typical displacement, we note that the impurity exhibits a diffusive motion, albeit with a decreasing self-diffusion coefficient. According to kinetic theory, the self-diffusion coefficient in the hard-sphere gas is the product of the mean-free path by thermal velocity Resibois and Leener (1977); Krapivsky et al. (2010); Dorfman et al. (2021). Thus, $D_{0}=\lambda_{0}\sqrt{T_{0}}=\lambda_{0}V_{\text{imp}}$. We use the mean-free path $\lambda_{0}=\lambda n_{\infty}/n_{0}$ inside the core, so the self-diffusion coefficient reads
$D_{0}=\lambda\,\frac{n_{\infty}}{n_{0}}\,V_{\text{imp}}=\lambda V\phi^{\frac{d-1}{2}}\,\left(\frac{\lambda}{R}\right)^{3\omega_{d}-d}$ (20)
The typical displacement of the impurity is found from $R_{\text{imp}}=\sqrt{D_{0}t}$ which in conjunction with Eqs. (20), (2), and (1) lead to the announced estimate (4).
Thus assuming that the impurity stays within the core region we have demonstrated the consistency by showing that the typical displacement of the impurity grows as the size of the core region. The impurity certainly stays behind the shock wave, but assuming that it explores the $r\leq R(t)$ ball and repeating the above line of reasoning gives $R_{\text{imp}}$ that grows slower than $R\propto t^{2/(d+2)}$. This failure has led us to the $R_{\text{imp}}\sim H$ guess that turned out to be self-consistent.
The hydrodynamic flow in the vicinity of impurity is radial, $u_{\text{imp}}\sim R_{\text{imp}}/t$. Thus,
$u_{\text{imp}}=V\phi^{\frac{d-1}{2}}\left(\frac{\lambda}{R}\right)^{\frac{d}{2}+1-h_{d}}$ (21)
Focusing on the time dependence, we have
$R_{\text{imp}}$ $\propto t^{\frac{2}{d+2}-\frac{8}{(d+2)(8+3d^{2})}}$ (22a)
$V_{\text{imp}}$ $\propto t^{-\frac{d}{d+2}+\frac{2d^{2}}{(d+2)(8+3d^{2})}}$ (22b)
$u_{\text{imp}}$ $\propto t^{-\frac{d}{d+2}-\frac{8}{(d+2)(8+3d^{2})}}$ (22c)
Note that $u_{\text{imp}}(t)$ decays faster than $V_{\text{imp}}(t)$. This justifies using $V_{\text{imp}}$ and neglecting the hydrodynamic radial velocity in estimating the self-diffusion coefficient.
To determine the number of collisions experienced by the impurity, we use the Boltzmann cylinder argument Resibois and Leener (1977); Dorfman et al. (2021) and estimate the time interval $\tau=(n_{0}a^{d-1}V_{\text{imp}})^{-1}$ between collisions at time $t$. The total number of collisions is estimated from $C_{\text{imp}}=t/\tau$. Using Eq. (7) and Eq. (18), we obtain
$C_{\text{imp}}=\left(\frac{R}{\lambda}\right)^{\frac{8+2d^{2}}{8+3d^{2}}}\propto t^{\frac{4(4+d^{2})}{(2+d)(8+3d^{2})}}$ (23)
Table 1 summarizes the scaling laws in two and three dimensions for the gas of hard spheres. Only the dependence on time is shown.
 $d$  $C_{\text{imp}}$  $R_{\text{imp}}$ $V_{\text{imp}}$ $u_{\text{imp}}$ $n_{0}$ $T_{0}$
2 $t^{2/5}$ $t^{2/5}$ $t^{-2/5}$ $t^{-3/5}$ $t^{-1/5}$ $t^{-4/5}$
3 $t^{52/175}$ $t^{62/175}$ $t^{-87/175}$ $t^{-113/175}$  $t^{-36/175}$ $t^{-174/175}$
Table 1: The temporal growth law for the number of collisions experienced by the impurity, Eq. (9), and the typical displacement of the impurity, Eq. (4); the latter quantity grows similarly to the size of the core region, Eq. (3). Two following columns present the decay laws for the typical speed of the impurity, Eq. (7), and the hydrodynamic radial velocity in the vicinity of the impurity, Eq. (21). The two last columns give the density, Eq. (18), and the temperature, Eq. (19), at the center of the explosion. We show the asymptotic behaviors in the physically relevant cases of two and three dimensions.
IV Position and momentum distributions
The collision dynamics is deterministic, but the outcome depends on the initial locations of the static spheres. Hence, we can treat the position and velocity of the impurity as random variables. The corresponding distributions $P(\bm{r},t)$ and $Q(\bm{v},t)$ cannot be separately studied since the position and velocity are correlated. For instance, if the impurity has a larger than average speed, it is expected to be further away from its initial position. Thus, only the joint distribution $\Pi(\bm{r},\bm{v},t)$ provides an adequate (and comprehensive) description. The joint distribution encodes the individual distributions:
[…]
Let us ignore correlations and try to probe the position distribution using a physically appealing but uncontrolled approximation. Employing a diffusion equation with a time-dependent diffusion coefficient is a naive approach. Indeed, the hydrodynamic radial velocity leads to $H\sim R_{\text{imp}}$ drift comparable with displacement generated by diffusion. A more consistent but still uncontrolled approximation is based on a convection-diffusion equation. Focusing on time dependence, ignoring numerical factors, and taking into account the spherical symmetry, we get a convection-diffusion equation
$\big{(}\partial_{t}+t^{-\frac{113}{175}}\partial_{r}\big{)}P=t^{-\frac{51}{175}}\big{(}\partial_{r}^{2}+\tfrac{2}{r}\partial_{r}\big{)}P$ (24)
in three dimensions. We have also taken into account the spherical symmetry. We seek the solution of the convection-diffusion equation (24) in the scaling form
$P(r,t)=t^{-\frac{186}{175}}\,\mathcal{P}(\eta),\qquad\eta=t^{-\frac{62}{175}}r$ (25)
The normalization requirement becomes
[…]
Plugging (25) into (24) we obtain a second-order ordinary differential equation
$-\frac{186}{175}\,\mathcal{P}-\frac{62}{175}\,\eta\,\mathcal{P}^{\prime}=\mathcal{P}^{\prime\prime}+\frac{2}{\eta}\,\mathcal{P}^{\prime}$ (27)
where prime denotes the derivative with respect to $\eta$. Multiplying Eq. (27) by $\eta^{2}$ and integrating yields the first-order ordinary differential equation
$-\frac{62}{175}\,\eta^{3}\mathcal{P}=\eta^{2}\mathcal{P}^{\prime}$ (28)
The solution to Eq. (28) satisfying the normalization requirement (26) is
$\mathcal{P}=\left(\frac{31}{175\pi}\right)^{\frac{3}{2}}\exp\!\left[-\frac{31}{175}\,\eta^{2}\right]$ (29)
The omitted numerical factors do not affect the Gaussian form. Thus, the convection-diffusion approach leads to the Gaussian solution in three dimensions. The extension to any $d\geq 2$ is straightforward and gives the same result:
$P(\bm{r},t)=\frac{1}{(2\pi R_{\text{imp}}^{2})^{d/2}}\,\exp\!\left[-\frac{r^{2}}{2R_{\text{imp}}^{2}}\right]$ (30)
with $R_{\text{imp}}^{2}=C_{d}\lambda^{2}(R/\lambda)^{2h_{d}}$. The numerical factor $C_{d}$ cannot be fixed in the realm of the above approach. We emphasize that have not justified the convection-diffusion approach, so the Gaussian form (30) of the position distribution is questionable.
A closed-form equation governing the evolution of the velocity distribution $Q(\bm{v},t)$, even a description based on uncontrolled approximation like the convection-diffusion equation for the position distribution, is unknown. The simplest guess is that the velocity distribution is asymptotically Gaussian. We do not have rational arguments in favor of this conjecture.
In contrast to the position and velocity distributions, the joint distribution satisfies a closed Lorentz-Boltzmann equation. In three dimensions
$(\partial_{t}+\bm{v}\cdot\nabla)\Pi(\bm{r},\bm{v},t)=4\pi a^{2}[\Gamma_{+}-\Gamma_{-}]$ (31a)
The terms on the right-hand side of Eq. (31a) account for collisions of the impurity with host spheres:
[…]
In the gain term (31b), we denote by $\bm{v}^{*}$ the pre-collision velocity of the impurity and by $\bm{w}^{*}$ the pre-collision velocity of a host sphere colliding with impurity
$\bm{v}^{*}=\bm{v}+\bm{e}[\bm{e}\cdot(\bm{v}-\bm{w})],\quad\bm{w}^{*}=\bm{w}-\bm{e}[\bm{e}\cdot(\bm{v}-\bm{w})]$ (32)
[…]
$F(\bm{r},\bm{w},t)=\frac{\rho(r,t)}{[2\pi T(r,t)]^{3/2}}\,\exp\!\left\{-\frac{m|\bm{w}-u(r,t)|^{2}}{2T(r,t)}\right\}$
The density $\rho(r,t)$, temperature $T(r,t)$, and radial velocity $u(r,t)$ satisfy the Navier-Stokes equations in the core region. In the long time limit, these scalar fields acquire the scaling form depending on a single variable $r/H$. The scaled density, temperature, and radial velocity satisfy non-linear coupled ordinary differential equations that have not been solved in the core region.
The governing Eq. (31) is a linear integro-differential equation that seems analytically intractable. Indeed, the distribution $F(\bm{r},\bm{w},t)$ describing host spheres contains the density, temperature, and radial velocity. Analytical expressions of these fields inside the core region are unknown, so the integral in Eq. (31c) is already unknown.
Another challenge is to analyze the number of collisions experienced by the impurity beyond the typical value, Eq. (9), estimated via arguments based on kinetic theory. Even for the dilute hard-sphere gas in thermal equilibrium, the collisional statistics of the tagged particle is remarkably subtle Lue (2005); Visco et al. (2008a, b); Paik (2014).
V Numerical verification
To test the predictions about the characteristics of impurities, we simulated hard discs in two dimensions using an event-driven molecular dynamics (EDMD) scheme. The system is prepared by randomly distributing discs, while ensuring the no overlap condition, within a simulation box of dimensions $L\times L$, at a fixed specified number density $n_{\infty}=N/L^{2}$ and volume fraction ($\phi=\pi a^{2}n_{\infty}$). We chose four discs closest to the origin and gave them random momenta, ensuring that the total energy is exactly $E$ and the total momentum vanishes. The simulation proceeds by calculating the times of all possible future collisions between discs and with the walls, then evolving the system to the earliest collision event. At each event, the velocities of the colliding particles are updated based on the conservation of momentum and energy. Between collisions, the discs move linearly with constant velocity. In all simulations performed, we used density $n_{\infty}=0.15$, the radius of particles $a=1/2$, and the mass of particles $m=1$.
The system contained $N=40000$ particles, and the EDMD simulations ran until $t=4000$. For computing the properties of impurities, we averaged over $30000$ randomly chosen initial conditions, while for computing bulk and surface quantities, we used $5000$ initial conditions. For any observable, say $A$, we expect power law behavior, $A\sim t^{\alpha}$, in the long time limit, while at finite times, there will be corrections. Hence, in our numerics, we estimate the temporal evolution of this exponent for various quantities of interest. We recorded observables at exponentially separated times in powers of $2^{1/4}$. We then obtain the exponents from the data at successive times (namely $\alpha(t)=4[\ln A(2^{1/4}t)-\ln A(t)]/\ln 2$).
$\bm{R}_{imp}$ $\bm{C}_{imp}$ $\bm{V}_{imp}$ $\bm{u}_{imp}$ $\bm{T}_{0}$ $\bm{n}_{0}$
$\alpha_{T}$ $2/5$ $2/5$ $-2/5$ $-3/5$ $-4/5$ $-1/5$
$\alpha_{S}$ $0.423$ $0.363$ $-0.399$ $-0.446$ $-0.692$ $-0.168$
$\delta\alpha_{S}$ $0.014$ $0.007$ $0.017$ $0.042$ $0.035$ $0.034$
Table 2: The table shows the comparison between theoretically computed and numerically estimated exponents $\alpha$ for time $(t)$ dependencies of the form $t^{\alpha}$ for the quantities of interests. $\alpha_{T}$ and $\alpha_{S}$ stands for Theory and Simulation.
Impurity properties: Let ${\bf r}(t)$ be the displacement, at time $t$, of any of the four impurity particles from its initial position and ${\bf v}(t)$ be its velocity. We define the typical impurity displacement and velocity as:
$R_{\rm imp}(t)=\langle{\bf r}^{2}/2\rangle^{1/2},~{}~{}~{}~{}V_{\rm imp}(t)=\langle{\bf v}^{2}/2\rangle^{1/2},$ (33)
where the average is over $n_{R}=3\times 10^{4}$ different initial conditions and over the four impurities. We define $C_{imp}(t)$ as the mean number of collisions that any impurity undergoes up to time $t$. The comparison between the kinetic predictions and the simulation results are shown in Fig. 2$(A)$-$(C)$.
To compute the fluid speed, $u_{imp}$, around the impurity particle, we find the average flow velocity of all the particles within a fixed radius around the impurity and compute its magnitude. The fixed radius is chosen so that, at $t=0$, there are $100$ other gas particles within that distance. Again, we take an average over the four impurities and $n_{R}$ realizations. The comparison of $u_{\text{imp}}$ between theory and simulations is shown in Fig. 2$(D)$.
In Fig. 3 we plot the full probability distribution function of the impurity’s displacement at time $t$. From Eq. 30, the predicted form is
$p(r,t)=2\pi rP({\bf r},t)=\frac{r}{R_{\rm imp}^{2}}\exp\left(-\frac{r^{2}}{2R_{\rm imp}^{2}}\right).$ (34)
Our simulations find close agreement with this.
Figure 3: Scaled probability distribution function for impurity displacement plotted at different times. The dashed curve in the figure is plotted by fitting Eq. (34) with $R_{\text{imp}}=b~{}t^{2/5}$ against the PDF obtained from simulation data at $t=4000$. The parameter $b$ obtained by fitting is $1.75$ which is close to $b=1.71$, obtained directly from computation of $R_{imp}/t^{2/5}$ at $t=4000$.
Core properties: For the computation of the quantities in the blast core, namely central density $n_{0}$ and central temperature $T_{0}$, we consider a circular region that includes $100$ particles at $t=0$. Then, we count the total energy and the number of particles in this circular region. Dividing the energy by the number of particles inside the circular region gives the central temperature, $T_{0}$, while dividing the number of particles by the area of the central region gives the density, $n_{0}$. The results for $n_{0}$ and $T_{0}$ are shown in Fig. 2(E)-(F).
A summary of our results for various exponent values and their comparison with the analytic prediction reported is given in Table (2). The reported exponent values correspond to the last two recorded times.
VI Summary
The large-scale motion generated by kicking an impurity in a homogeneous cold gas is equivalent to the hydrodynamic behavior in the earlier stage of the atomic explosion when the pressure behind the shock wave greatly exceeds the atmospheric pressure. The hydrodynamic behavior is tractable in this regime. The pressure ahead of the shock vanishes in the zero-temperature gas, so the shock remains infinitely strong, and the analytical solution applies forever.
We analyzed the evolution of the impurity. We probed the typical displacement, the average speed, and the average number of collisions experienced by the impurity. The chief observation is that the impurity remains inside the core region where the standard description of the infinitely strong blast based on the Euler equations becomes inaccurate, and one should rely on the Navier-Stokes equations accounting for dissipative processes.
The impurity in the cold homogeneous hard-sphere gas exhibits algebraic scaling behaviors with amusing rational exponents (Table 1). Our derivations of these scaling laws rely on arguments based on elementary kinetic theory along with inputs from hydrodynamics both at the Euler and dissipative scales. In dimension $d=2$, we performed molecular dynamics simulations to test our heuristic predictions and find reasonable agreement. We believe that simulations at longer times will lead to better agreement.
The linear Lorentz-Boltzmann equation describing the evolution of the joint position-velocity distribution of the impurity provides the theoretical framework for the comprehensive analysis of the impurity. Unfortunately, the Lorentz-Boltzmann equation appears analytically intractable. In addition to the mathematical challenges caused by the necessity to solve an integro-differential equation in an inhomogeneous and evolving background, we do not even have the analytical description of the background due to the lack of the analytical solution of the Navier-Stokes equations in the core region.
Impurities are generally different from the host particles. The disparity between the mass and radius of impurity and the mass and radii of host spheres may lead to interesting behaviors, especially when these disparities are significant. We leave this problem for the future. One can also explore the impurity problem in high spatial dimensions, $d\gg 1$, which may be more amenable to analytical treatment and perhaps sheds light on the behavior in the physically relevant three-dimensional case. Finally, we mention a different setting where the impurity is normally incident on the static gas occupying the half-space Antal et al. (2008). The impurity, as well as every host sphere, is expected to be ejected backward into an initially empty half-space after a finite number of collisions. This assertion is intuitively clear, albeit proof could be challenging. So far, the splash problem has been quantitatively studied in one-dimension Chakraborti et al. (2022); Kumar and Rajesh (2025b).
VII Acknowledgement
We thank Subhadip Chakraborti and Frank Smallenburg for useful discussions on numerical simulation techniques. AD and UK acknowledge support from the Department of Atomic Energy, Government of India, under project No. RTI4001. AD acknowledges the J.C. Bose Fellowship (JCB/2022/000014) of the Science and Engineering Research Board of the Department of Science and Technology, Government of India.